Suppose a country produces two final goods only, good A and good B. In 2017, the price for good A was $3 and the price for good B was $7. In 2018, the prices for good A and good B were $3 and $11 respectively, and the outputs of good A and B were 103 units and 202 units respectively. In 2019, the prices for good A and good B are $7 and $11 respectively, the outputs are 122 units and 238 units respectively. The base year is 2017. What is the growth rate of real GDP from 2018 to 2019?

Round your answer to 2 digits after the decimal point. Leave out the percentage sign, that is, if your answer is 1.23%, type 1.23, not 1.23%, not 0.0123.
Answer #1

Ans: The growth rate of real GDP from 2018 to 2019 is 17.93

Explanation:

Real GDP = Sum of base year price * Current year quantity

Real GDP for 2018 : ( 3 * 103 ) + ( 7 * 202 ) = 309 + 1414 = 1723

Real GDP for 2019 : ( 3 * 122 ) + ( 7 * 238 ) = 366 + 1666 = 2032

The growth rate of real GDP from 2018 to 2019 = {(Real GDP for 2019 - Real GDP for 2018) / Real GDP for 2018} *100

= {( 2032 - 1723 ) / 1723} *100 = (309 / 1723 ) *100 = 0.1793 * 100 = 17.93